# ConceptWM: Watermarking Diffusion Models for Visual Concept Copyright Protection

### v2

@@ 4.1. Implementation Details
Concept-Oriented Ours(OLW-based) 24.06 0.357 98.68 90.03 0.998 0.884
Ours(FLW-based) 24.58 0.365 95.68 89.32 0.994 0.858
Table 1. A comparison between our method and previous watermarking techniques is presented. The capacities for DWT-DCT-SVD, RivaGAN, StableSignature, and our method are 48 bits, 32 bits, 48 bits, and 48 bits, respectively. We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR). Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case. ”Adv.” (Adversarial) refers to the average performance of images under various distortions.
Distortion Layer during training. During training, several data augmentation techniques were employed to enhance the model’s robustness. We follow the setup of attack layers in the representative work Stable Signature (Fernandez et al., 2023). Specifically, JPEG compression was simulated by using the HiDDeN (Zhu et al., 2018) model’s JPEG distortion layer, with compression quality randomly chosen between 40 and 100 to simulate different levels of image compression. Random cropping was applied, resizing the cropped image within the range of [256, 512] pixels, and then scaling it back to 512×512 pixels to ensure variability in the region of interest. Gaussian blur was applied with kernel sizes randomly selected between [3, 9] and blur intensities ranging from 0 to 2, simulating distortions from defocus or motion blur. Gaussian noise with a mean of 0 and variance of 10 was added to simulate low-quality or noisy images. Color jittering adjusted brightness, contrast, and saturation within the range of [0.8, 1.25], and hue was randomly shifted between [-0.2, 0.2] to simulate lighting and color variations. These augmentations effectively improved the model’s adaptability to compression, noise, blur, and color variations, ensuring stable watermarking even under diverse image transformations.
Evaluation Metrics. We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5. For traceability performance, we assess bit accuracy. To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou, 2010) and Structural Similarity Index (SSIM) (Wang et al., 2004) to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) (Mittal et al., 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al., 2015) for evaluating overall image quality.
Watermarking baselines. We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al., 2022) called DwtDctSvd (Cox et al., 2007), and RivaGAN (Zhang et al., 2019),VAE-based method Stable Signature (Fernandez et al., 2023) and Latent-based method Tree-ring (Wen et al., 2024). DwtDctSvd, RivaGAN, and Stable Signature all use 48-bit watermarks, while Tree-ring uses P-value as the evaluation method.
@@
Contrast Brightness Blur Noise Jpeg Crop&Rescale Vae.C Denoising Average
Bit Accuracy(%$\uparrow$ Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
@@
Ours (FLW-based) 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
Table 2. The performance of ConceptWM against various types of attacks is compared with other watermarking schemes. The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023c).
4.2. Watermark Performance Comparison
We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) is set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images. (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks. Specifically, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing. Each user generates 10 images, creating a dataset of 10,000 watermarked images. During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images generated by the user. The user with the highest Bit Accuracy is considered the traced user.
@@ 4.2. Watermark Performance Comparison
4.3. Fidelity
Table 1 presents a comparison of our method with other baseline methods. For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al., 2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al., 2014). Additionally, we use CLIP score (Radford et al., 2021), a method for measuring image similarity, which provides results more aligned with human judgment.
Settings BitAcc(%) TPR (Fpr=$10^{-6}$)
@@ 4.3. Fidelity
Table 3. The impact of inference settings on ConceptWM watermark detection.
Main Results. The performance of this method outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02). This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance. Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods. It can also be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower. In contrast, the image quality produced by FLW, has an (FID) of 24.58 and semantic quality (CLIP) of 0.365. This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark. Ultimately the latter leads to a degradation in image quality.
4.4. Watermark Robustness against Attack
Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023d). These cover common image processing techniques, geometric transformations, and regeneration attacks encountered in real-world scenarios. The main results are shown in Table 2.
@@ 4.4. Watermark Robustness against Attack
Table 4. An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization. The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
@@
Table 5. The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
4.5. Concept Protection Evaluation
@@
Textual Inversion and DreamBooth with LoRA. Textual Inversion teaches new concepts by adjusting a word vector, rather than fine-tuning the entire model. LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency and is widely adopted within the community. The relevant results are shown in Table 4. Our method shows better performance in mitigating the degradation effect on generation with LoRA. Since Textual Inversion does not inject concepts directly into the model, the reduction in quality is not significant and the semantic quality of the generation decrease.
@@
Prompt Mismatching. Considering that users generate relevant images based on various prompts combined with specific concepts, we generate images for specific concepts under different prompt conditions. The main results are shown in Table 4, which shows that our watermarking method achieves higher bit extraction accuracy for images similar to the training set. When the image deviates significantly from the set instance prompt, our watermarking method still maintains good precision. Furthermore, the generated watermark exhibits a certain level of adversarial robustness, as the quality of images trained under this watermark is reduced, which is beneficial for the protection of specific concepts in diffusion models.
@@
Distortion Layer. We evaluated the impact of various attacks on image quality in the Distortion Layer in the watermark components pretraining. The main results are in Table 5, which indicates that cropping and the mask adversarial layers have a significant impact on image quality, as they require a stronger watermark to be stably embedded within a smaller space.
@@
Inference Steps. Users may influence watermark extraction under different inference parameters. We generated images of specific concepts at different inference steps and performed watermark extraction. The results, shown in Table 3, indicate that the ConceptWM remains effective even with varying inference steps.
Guidance Scales. Larger guidance scales result in more faithful of the generated image adherence to prompts. We evaluate the impace of guidance scale in the Table 3. The results show that our method maintains good stability across various values of the Guidance scale.
Sampling Method. We evaluate our watermark on DDIM (Song et al., 2020), DPM-Solver multi-step (Lu et al., 2022), Euler and Heun samplers (Karras et al., 2022). The results show that our watermark is minimally affected by the choice of sampler.
@@
Different VAE Decoder. An adversary may interfere with watermark decoding by replacing the VAE decoder. To evaluate the impact of the generated images on the watermark, we used different types of VAEs including stable diffusion V1.4 and V2.1 for decoding. The results in Table 3 indicate that changing the VAE does not have a significant impact on the watermark extraction, which is a clear advantage over VAE-based schemes, as the watermark cannot be preserved when the VAE is replaced in VAE-based methods.
@@
Generation Size. The results in Table 3 show that different generation sizes have a noticeable impact on watermark, while the watermarking method still remains effective and detectable across different generation sizes.
5. Conclusion
@@
This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts. Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results. We have evaluated the stability and adversarial effectiveness of the watermarking method we designed, demonstrating that our defense remains effective under adverse conditions. In the future, our main focus will be on two aspects of improvement. On the one hand, we aim to address the efficiency issue caused by the need to generate adversarial noise in the current method. On the other hand, we will consider potential threats from attackers who may target the model level.

### v3

@@ Introduction
Concept-Oriented OLW-based 24.06 0.357 98.68 90.03 0.998 0.884
FLW-based 24.58 0.365 95.68 89.32 0.994 0.858
Table 1: A comparison between our method and previous watermarking techniques is presented. We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR). Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case. “Adv.” (Adversarial) refers to the average performance of images under various distortions.
Experiment
Datasets. To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al. 2022). For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period. Each image is resized and center-cropped to a resolution of 512 × 512. In each experiment, a total of 5–8 carefully selected images were utilized for DreamBooth fine-tuning.
[…]
@@
Contrast Brightness Blur Noise Jpeg C.&R. Vae.C Denoising Average
Bit Accuracy(%) Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
@@
Ours 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
Table 2: The performance of ConceptWM against various types of attacks is compared with other watermarking schemes. The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al. 2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al. 2023a).
Evaluation Metrics. We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5. For traceability performance, we assess bit accuracy. To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou 2010) and Structural Similarity Index (SSIM) (Wang et al. 2004); and the Natural Image Quality Evaluator (NIQE) (Mittal, Soundararajan, and Bovik 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al. 2015) for evaluating overall image quality.
Watermarking baselines. We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al. 2022) called DwtDctSvd (Cox et al. 2007), and RivaGAN (Zhang et al. 2019),VAE-based method Stable Signature (Fernandez et al. 2023) and Latent-based method Tree-ring (Wen et al. 2024). DwtDctSvd, RivaGAN, and Stable Signature all use 48-bit watermarks, while Tree-ring uses P-value as the evaluation method.
Watermark Performance Comparison
We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images. (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks. Specifically, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing. Each user generates 10 images, creating a dataset of 10,000 watermarked images. During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images generated by the user.
@@
Watermarked Diffusion Model Fidelity
For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al. 2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al. 2014). Additionally, we use CLIP score (Radford et al. 2021), a method for measuring image similarity, which provides results more aligned with human judgment. Table 1 presents a comparison of our method with other baseline methods.
Settings Bit Accuracy TPR
@@
Table 3: The impact of inference settings on ConceptWM watermark detection.
Main Results. The performance of our method outperforms the state-of-the-art methods, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02). This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance. Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods. It can be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower. In contrast, the image quality produced by FLW, has an FID of 24.58 and semantic quality (CLIP) of 0.365. This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark, which leads to degradation in image quality.
Watermark Robustness against Attack
Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al. 2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al. 2023b). These cover common image processing techniques, geometric transformations, and regeneration attacks encountered in real-world scenarios. The main results are shown in Table 2.
@@
Table 4: An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization. The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
@@
Table 5: The impact of various attacks on image quality in the distortion layer during pretraining.
Concept Protection Evaluation
@@
Textual Inversion and DreamBooth with LoRA. Textual Inversion teaches new concepts by learning a new token, rather than fine-tuning the entire model. LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency. The relevant results are shown in Table 4. Our method shows better performance in mitigating the degradation effect on generation with LoRA. Since Textual Inversion does not inject concepts directly into the model, the reduction in quality is not significant and the semantic quality of the generation decrease.
@@
Prompt Mismatching. Considering that users generate relevant images based on various prompts combined with specific concepts, we generate images for specific concepts under different prompt conditions. The results are shown in Table 3. When the image deviates significantly from the set instance prompt, our watermarking method still maintains bit accuacy higher than 0.8.
@@
Distortion Layer. We evaluate the impact of Distortion Layer in the watermark components pretraining. The main results are in Table 5, which indicates that cropping and the mask adversarial layers have a significant impact on image quality, as they require a stronger watermark to be stably embedded within a smaller space.
@@
Inference Steps. As shown in Table 3, the bit accuracy remains stable at approximately 0.93 across different settings - specifically reaching 0.917 at 25 steps, 0.936 at 50 steps, and 0.936 at 100 steps. The results indicate that the ConceptWM remains effective with varying inference steps.
Guidance Scales. Our experiments demonstrate that ConceptWM maintains robust watermark detection performance across varying guidance scales, achieving 0.932 bit accuracy at scale 5, 0.917 at scale 7.5, and 0.933 at scale 10.
Sampling Method. Evaluation on DDIM (Song, Meng, and Ermon 2020), DPM-Solver (Lu et al. 2022), Euler and Heun samplers (Karras et al. 2022) shows consistent performance with 0.929, 0.938, 0.922 and 0.947 bit accuracy respectively, demonstrating minimal impact from sampler choice.
@@
Different VAE Decoder. An adversary may interfere with watermark decoding by replacing the VAE decoder. We used different types of VAEs including stable diffusion V1.4 and V2.1 for decoding, achieving 0.901 and 0.939 bit accuracy respectively. The results in Table 3 indicate that changing the VAE does not have a significant impact on the watermark extraction, which is a clear advantage over VAE-based schemes, as the watermark cannot be preserved when the VAE is replaced in VAE-based methods.
@@
Generation Size. The results in Table 3 show that different generation sizes have some impact on watermark accuracy, with 0.933 for 512×512 and 0.812 for 768×768 generations, while the watermarking method still remains effective and detectable across different generation sizes.
Conclusion
@@
This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts. Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results. In the future, The investigation of concept-specific learning can be extended beyond DreamBooth, necessitating an in-depth examination of watermark embedding paradigms for textual inversion, LoRA, and related concept learning methodologies.